# Federated Survival Analysis with Site-Level Heterogeneity Adjustment

### 10 Mar 2026 v2

@@ 2.1 Notation
We begin by defining key notations in survival analysis. Let $T_{i}$ and $C_{i}$ denote the true event time and censoring time, respectively, for subject $i=1,\dots,N$. We observe the follow-up time $X_{i}=\min(T_{i},C_{i})$ and the event indicator $\Delta_{i}=\mathbb{I}(T_{i}\leq C_{i})$, where $\mathbb{I}(\cdot)$ is the indicator function. Let $Z_{i}$ denote a $p$-dimensional vector of baseline covariates for subject $i$. The observed data for the cohort is denoted by $\mathcal{D}=\{(X_{i},\Delta_{i},Z_{i})\}_{i=1}^{N}$.
@@ 2.2 Review Pseudo-value Regression in a Single Site
where $t_{1}<t_{2}<\cdots$ denote the ordered distinct event times, $d_{k}$ is the number of events at time $t_{k}$, and $Y_{k}$ is the number of individuals at risk immediately prior to $t_{k}$. Based on this estimator, the pseudo-observation (Andersen and Pohar Perme, 2010) for subject $i$ evaluated at a prespecified landmark time $t_{j}$ is constructed via a leave-one-out (jackknife) approach, which is defined as $\tilde{S}_{ij}=N\hat{S}(t_{j})-(N-1)\hat{S}^{(-i)}(t_{j}),$ where $\hat{S}(t_{j})$ is the KM estimator computed using all $N$ subjects and $\hat{S}^{(-i)}(t_{j})$ is the KM estimator computed after removing subject $i$ from the sample. These pseudo-observation $\tilde{S}_{ij}$ are treated as continuous outcomes and modeled using a generalized linear model (GLM) to estimate covariate effects $Z_{i}$ on the survival probabilities at time $t_{j}$ :
@@
To model survival probabilities over the follow-up period, pseudo-observations are typically evaluated at multiple landmark times $t_{1},\dots,t_{J}$ (Andersen and Pohar Perme, 2010; Anyaso-Samuel et al., 2023; Zhao and Feng, 2020; Zhao, 2020). It has been shown that as few as five time intervals equally spread on the event time scale worked quite well in most cases (Zhao and Feng, 2020; Zhao, 2020). Stacking these pseudo-observations yields repeated outcomes for each subject. Let $\theta_{i}=(\tilde{S}_{i1},\dots,\tilde{S}_{iJ})^{\top}$ denote the vector of pseudo-observations for subject $i$. Regression parameters are estimated by solving the generalized estimating equations (GEE):
@@ 3.2 Scenario 2: Time-Varying Treatment Effects
To evaluate model performance when the PH assumption is violated, we generated survival data from a Weibull model incorporating a binary treatment indicator $A$ and a continuous confounder $X$. The shape parameter was defined as a function of treatment, $\alpha(A)=\alpha_{0}+\gamma A$, while the scale parameter followed $\lambda(X,A)=\lambda_{0}\exp(\beta_{X}X+\beta_{A}A)$. This specification induces a time-dependent HR, thereby violating the PH assumption when $\gamma\neq 0$.
We simulated N=5,500N=5{,}500 subjects distributed across $K=20$ sites, with site-specific sample sizes ranging from 50 to 3,000. For each subject, we simulated a continuous confounder $X\sim\mathcal{N}(0,1)$ and assigned a binary treatment $A$ via $\text{logit}\{P(A=1\mid X)\}=0.5X$. Parameters were set to $\alpha_{0}=0.8$, $\gamma=0.5$, $\lambda_{0}=15$, $\beta_{X}=\log(0.7)$, and $\beta_{A}=\log(1.5)$.

### 11 Mar 2026 v3

@@ 2.1 Notation
We begin by defining key notations in survival analysis. Let $T_{i}$ and $C_{i}$ denote the true event time and censoring time, respectively, for subject $i=1,\dots,N$. We observe the follow-up time $X_{i}=\min(T_{i},C_{i})$ and the event indicator $\Delta_{i}=\mathbb{I}(T_{i}\leq C_{i})$, where $\mathbb{I}(\cdot)$ is the indicator function. Let $Z_{i}$ denote a $p$-dimensional vector of baseline covariates for subject ii. The observed data for the cohort is denoted by $\mathcal{D}=\{(X_{i},\Delta_{i},Z_{i})\}_{i=1}^{N}$.
@@ 2.2 Review Pseudo-value Regression in a Single Site
where $t_{1}<t_{2}<\cdots$ denote the ordered distinct event times, $d_{k}$ is the number of events at time $t_{k}$, and $Y_{k}$ is the number of individuals at risk immediately prior to $t_{k}$. Based on this estimator, the pseudo-observation (Andersen and Pohar Perme, 2010) for subject $i$ evaluated at a prespecified landmark time $t_{j}$ is constructed via a leave-one-out (jackknife) approach, which is defined as $\tilde{S}_{ij}=N\hat{S}(t_{j})-(N-1)\hat{S}^{(-i)}(t_{j}),$ where $\hat{S}(t_{j})$ is the KM estimator computed using all NN subjects and $\hat{S}^{(-i)}(t_{j})$ is the KM estimator computed after removing subject $i$ from the sample. These pseudo-observation $\tilde{S}_{ij}$ are treated as continuous outcomes and modeled using a generalized linear model (GLM) to estimate covariate effects $Z_{i}$ on the survival probabilities at time $t_{j}$ :
@@
To model survival probabilities over the follow-up period, pseudo-observations are typically evaluated at multiple landmark times $t_{1},\dots,t_{J}$ (Andersen and Pohar Perme, 2010; Anyaso-Samuel et al., 2023; Zhao and Feng, 2020; Zhao, 2020). It has been shown that as few as five time intervals equally spread on the event time scale worked quite well in most cases (Zhao and Feng, 2020; Zhao, 2020). Stacking these pseudo-observations yields repeated outcomes for each subject. Let $\theta_{i}=(\tilde{S}_{i1},\dots,\tilde{S}_{iJ})^{\top}$ denote the vector of pseudo-observations for subject ii. Regression parameters are estimated by solving the generalized estimating equations (GEE):
@@ 3.2 Scenario 2: Time-Varying Treatment Effects
To evaluate model performance when the PH assumption is violated, we generated survival data from a Weibull model incorporating a binary treatment indicator AA and a continuous confounder $X$. The shape parameter was defined as a function of treatment, $\alpha(A)=\alpha_{0}+\gamma A$, while the scale parameter followed $\lambda(X,A)=\lambda_{0}\exp(\beta_{X}X+\beta_{A}A)$. This specification induces a time-dependent HR, thereby violating the PH assumption when $\gamma\neq 0$.
We simulated N=5,500N=5{,}500 subjects distributed across $K=20$ sites, with site-specific sample sizes ranging from 50 to 3,000. For each subject, we simulated a continuous confounder $X\sim\mathcal{N}(0,1)$ and assigned a binary treatment AA via $\text{logit}\{P(A=1\mid X)\}=0.5X$. Parameters were set to $\alpha_{0}=0.8$, $\gamma=0.5$, $\lambda_{0}=15$, $\beta_{X}=\log(0.7)$, and $\beta_{A}=\log(1.5)$.